OpenStudy (anonymous):

y=square root of 2x+3

OpenStudy (anonymous):

what about it?

OpenStudy (shamil98):

what are you looking for?

OpenStudy (anonymous):

how do u find the points of discontinuity and the type of discontinuity

OpenStudy (shamil98):

y = \[y = \sqrt{2x + 3}, y = \sqrt{2x} + 3\] which one is it?

OpenStudy (anonymous):

the former

OpenStudy (shamil98):

the points of discontinuity are found by finding out what x value makes the equation undefined.

OpenStudy (shamil98):

is there no fraction?

OpenStudy (anonymous):

no fraction, no

OpenStudy (shamil98):

hm..

OpenStudy (anonymous):

I know how to do most of like the other problems but with the square root I'm getting kind of thrown off

OpenStudy (shamil98):

There is no fraction, so i don't think there is any value that would make this undefined..

OpenStudy (anonymous):

all points in the domain i.e., all x<-3/2

OpenStudy (shamil98):

that's the global minimum

OpenStudy (shamil98):

as well

OpenStudy (shamil98):

but the fraction is just \[y = \frac{ \sqrt{2x+3} }{ 1 }\]

OpenStudy (anonymous):

so basically as long as the number in the root isn't less than 0 it's fine?

OpenStudy (shamil98):

Pretty much yes, you cant have something like √-25

OpenStudy (shamil98):

You only want real numbers ...

OpenStudy (shamil98):

\[inside \ge 0\]

OpenStudy (anonymous):

Is this a continuous function?

OpenStudy (shamil98):

I don't think so. The point of discontinuity would be (-3/2, 0) because all real numbers except at that value.

OpenStudy (anonymous):

ok thanks